A particular group of men have heights with a mean of 177 cm and a standard deviation of 7 cm. Charley had a height of 181 cm. a. What is the positive difference between Charley​'s height and the​ mean? b. How many standard deviations is that​ [the difference found in part​ (a)]? c. Convert Charley​'s height to a z score. d. If we consider​ "usual" heights to be those that convert to z scores between minus−2 and​ 2, is Charley​'s height usual or​ unusual?
Answer #1

(a)

difference= Charley height - mean

= 181-177 = 4cm

(b)

difference / Standard Deviation

= 4 / 7

(c)

z = (x-mean) / Standard Deviation

= (181-177) / 7 = 0.5714

(d)

This falls within the (-2, 2) range and the height would not be considered unusual.